OpenStudy (anonymous):

At the fair, Jamal's dad bought 5 hot dogs and 4 drinks for 14.75. Lana's mom bought 3 hot dogs and 6 drinks for 14.25. How much was one hot dog?

OpenStudy (anonymous):

haha

OpenStudy (gw2011):

Let H=Hot Dogs Let D=Drinks 5H+4D=14.75 3H+6D=14.25 Multiply the first equationby 6 and the second by 4 in order to find H and you get: 30H+24D=88.50 12H+24D=57 Subtract this equation from the first one and you get: 18H=31.50 H=$1.75 This is the price of one hot dog